# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ Abstract
Recent advancements in large language models (LLMs) have revolutionized their ability to handle single-turn tasks, yet real-world applications demand sophisticated multi-turn interactions. This survey provides a comprehensive review of recent advancements in evaluating and enhancing multi-turn interactions in LLMs. Focusing on task-specific scenarios—from instruction following in diverse domains such as math and coding to complex conversational engagements in roleplay, healthcare, education, and even adversarial jailbreak settings—we systematically examine the challenges of maintaining context, coherence, fairness, and responsiveness over prolonged dialogues. The paper organizes current benchmarks and datasets into coherent categories that reflect the evolving landscape of multi-turn dialogue evaluation. In addition, we review a range of enhancement methodologies under multi-turn settings, including model-centric strategies (contextual learning, supervised fine-tuning, reinforcement learning, and new architectures), external integration approaches (memory-augmented, retrieval-based methods, and knowledge graph), and agent-based techniques for collaborative interactions. Finally, we discuss open challenges and propose future directions for research to further advance the robustness and effectiveness of multi-turn interactions in LLMs. Related resources and papers are available at https://github.com/yubol-cmu/Awesome-Multi-Turn-LLMs.
footnotetext: Equal contribution.
@@ 1 Introduction
The advent of Large Language Models (LLMs), exemplified by influential systems such as GPT series [1, 2, 3], PaLM [4], and LLaMA [5], has significantly reshaped numerous domains, from education and healthcare to customer service and software engineering. These powerful language models demonstrate remarkable proficiency in generating coherent and contextually relevant responses, achieving groundbreaking performances across various language understanding and generation benchmarks.
However, much of the early progress in both the evaluation and improvements of LLMs has been concentrated in single-turn interactions, where models are tested on isolated prompts without considering prior conversational context. While this has led to impressive performance on a range of benchmarks, it overlooks the broader potential of LLMs in multi-turn dialogue, which more accurately reflects real-world usage. In practice, real-world scenarios rarely consist of isolated queries; instead, meaningful and productive interactions usually occur through continuous, multi-turn exchanges. Effective communication between humans and artificial intelligence inherently demands an understanding of conversational history, nuanced interpretation of previous exchanges, iterative refinement of goals, and adaptive response strategies. Single-turn interaction thus presents a significant limitation, restricting the deployment and utilization of the robust capabilities that LLMs possess.
@@
This emerging field presents rich opportunities and considerable challenges, prompting a rapidly expanding body of research dedicated to optimizing, evaluating, and deploying LLMs within multi-turn conversational settings. Understanding these multi-turn dynamics and systematically addressing their inherent complexities is essential for unlocking the next stage in LLM evolution, thereby significantly expanding their applicability and effectiveness in real-world scenarios.
@@
To provide clarity and facilitate further advancements, this survey specifically addresses multi-turn interaction with LLMs, categorized by tasks along with corresponding evaluation standards, improvement techniques, and current challenges in this domain. We categorize multi-turn interactions according to two main tasks: instruction following tasks and conversational engagement tasks, with the latter encompassing various real-world domain applications. While we refer to several LLM agent-based approaches, we emphasize that LLM-based agents represent a distinct research direction that, though related, falls outside our primary scope. Similarly, despite the growing body of research in Multi-modal LLMs (MLLMs) - including notable benchmarks and frameworks such as ConvBench [6], MMDU [7], MMMT-IF [8], MMDialog [9], TheaterGen [10], and SVBench [11] - this survey deliberately excludes multimodal capabilities to maintain a focused analysis of text-based multi-turn interactions. Such boundary allows us to provide a comprehensive exploration and in-depth analysis of the landscape of multi-turn interactions specific to LLMs.
@@
Foundational surveys [12, 13] provide rigorous analyses of LLMs’ ability to follow instructions under a single-turn setting. These works evaluate metrics like faithfulness, robustness, and generalization in static, one-off interactions. While they establish critical baselines for instruction alignment, their scope excludes the dynamic, context-dependent challenges inherent to multi-turn scenarios, such as coherence across turns, state tracking, and user intent adaptation.
Several dialogue system surveys are also particularly relevant: [14] present a overview of recent advances in multi-turn dialogue systems based on LLMs; healthcare-specific dialogue systems are examined by surveys [15, 16]. Additionally, [17] specifically explores multi-turn interaction capabilities of LLMs, with a strong focus on core model abilities, evaluation metrics, and algorithmic enhancements for multi-turn contexts.
@@
While existing instruction-following surveys predominantly address single-turn scenarios, and dialogue system surveys excel in targeted evaluations and performance considerations within their respective scopes, a clear gap emerges regarding a structured and holistic comparison under broader, multi-turn interaction settings.
Among current literature, Zhang et al. (2025)’s work [17] aligns most closely with our work. While it focuses primarily on LLM core multi-turn interaction capabilities, corresponding evaluation methods, and algorithmic improvements, their approach is fundamentally capability-oriented rather than task-oriented. In contrast, we recognize that multi-turn settings more closely resemble real-world applications, where performance emerges from the complex collaboration of multiple capabilities rather than isolated ones. Therefore, we deliberately categorize interactions by tasks instead of capabilities, offering a more practical taxonomy. Furthermore, Zhang et al. (2025) [17]’s survey lacks substantial discussion of multi-turn interactions in critical real-world scenarios, such as healthcare consultations and educational assistance, where multi-turn dynamics are essential. Their survey also presents limited analysis of improvement methodologies and future challenges in multi-turn contexts.
@@
To the best of our knowledge, this survey is the first to provide an extensive landscape of multi-turn LLM interactions, covering tasks, real-world applications, evaluation methods, improvement strategies, and critical open challenges.
Dataset Size Data Curation Evaluator Date***We use the date of their initial version on arXiv to rank them, allowing the development trajectory and patterns to be revealed. Benchmark/Dataset Category Total #Dial. Avg. #Turns/Dial. Human-based LLM-based Rule-based Human-as-a-judgeDo not include the human evaluation done for agreement check between LLM and human. LLM-as-a-judge Agreement CheckIf LLM-as-a-judge is used, whether there is an agreement check between LLM and human annotators on a subset of the dataset. Evaluation Criteria Mar 2022 MTPB [18] Coding 115 / no no yes no no / Pass rate, Perplexity Dec 2022 ABC-Eval §§§We intentionally include ABC-Eval and BotChat in this table as they represent early explorations in multi-turn interactions. Their presence illustrates the developmental trajectory of the field, although a more detailed discussion can be found in §2.2.  [19] Conversation 400 30.3 yes no no yes no / Consistency, Emotion, Understanding, Engagingness, Grammaticality, Informativeness, Quality, Proactivity, Relevance Jun 2023 MathChat-Agent [20] IF-math / 15 no no yes no no / Accuracy Jun 2023 MT-Bench [21] IF-general 80 2 yes no no no yes yes Correctness, Helpfulness Jun 2023 InterCode [22] IF-coding / 10 no yes yes no no / Execution Success Sep 2023 MINT [23] IF-reasoning 586 <5absent5<5< 5 no yes yes no no yes Evaluation Quality, Success Rate Oct 2023 MT-Bench++ [24] IF-general 80 8 no yes no no yes no Helpfulness, Relevance, Accuracy, Depth, Creativity, Level of Detail Oct 2023 BotChat [25] Conversation 547 16 yes yes no no yes / Quality, Similarity to Human Dialogues Jan 2024 EHRAgent [26] IF-coding / / yes no yes no no / Success Rate, Complete Rate Jan 2024 MT-Eval [27] IF-general 168 6.96 no yes yes no yes no Helpfulness, relevance, accuracy, depth, creativity, how well it conforms to constraints; accuracy of classification and recollection Jan 2024 WebLINX [28] IF-general 2337 43 yes no yes no no / Intent Match, Element Similarity Using IoU, Text Similarity Using F1, Turn-Level Score and Overall Score Feb 2024 AQA-Bench [29] IF-reasoning / 15,3015 for easy mode, 30 for hard mode. / / yes no no / Goal Metric, Policy Metric (Both for Each Interactive Environment) Feb 2024 MT-Bench-101 [30] IF-general 1388 3.03 no yes no no yes yes Perceptivity (Context Memory, Context Understanding, Context Interference), Adaptability (Rephrasing, Reflection, Reasoning), Interactivity (Questioning) May 2024 MathChat-Bench [31] IF-math 5276 4 no yes no no yes yes Follow-Up QA, Error Correction, Error Analysis, Problem Generation Jun 2024 M2Lingual M2Lingual includes dialogues for 70+ languages. The statistics shown here are restricted to only English multi-turn dialogues. [32] IF-general 1140 2.48 no yes MT-Bench same as MT-Bench’s Aug 2024 SysBench [33] IF-general 500 5 yes no yes no no / Constraint Satisfaction Rate, Instruction Satisfaction Rate, Session Stability Rate Oct 2024 FB-Bench [34] IF-general 591 2 yes yes no no yes yes Error Correction, Response Maintenance Oct 2024 WILT [35] IF-reasoning 50 <30absent30<30< 30 / / yes no no / Accuracy in Deducing the Hidden Functions Oct 2024 Multi-IF [36] IF-general 909 3 no yes yes no no / Instruction- and Conversation-Level Accuracy of Instruction Following Oct 2024 FairMT-Bench [37] IF-fairness 10k 4 no yes no no yes yes Direct and Implicit Bias Dec 2024 MMSQL [38] IF-coding 6493 6 no yes yes no yes yes Exact Matching, Execution Accuracy, Dual Assessment of Question Type Detection, Execution Accuracy, Response Quality Score
Table 1: An Overview of Recent Multi-Turn Instruction Following Bechmark or Datasets and Evaluation Methods.
@@ 2 Multi-Turn Interaction Tasks
In this survey, we categorize multi-turn interactions based on tasks rather than capabilities because real-world multi-turn scenarios inherently require integrating multiple capabilities to accomplish a particular objective. Although capabilities such as reasoning, memory, contextual understanding, or adaptability are critical, they rarely function in isolation within conversational contexts. Instead, different capabilities intertwine and complement each other dynamically to achieve specific task goals. By focusing on clearly defined tasks—such as multi-turn instruction-following or conversational engagement—we better reflect the realistic complexities of these interactions, offering readers an intuitive framework for understanding how LLMs practically meet user requirements and perform in real-world multi-turn conversational scenarios. Table 1 provides an overview of recent multi-turn instruction-following benchmarks or datasets and their evaluation methods, offering readers a detailed reference for understanding current trends and practices in multi-turn LLM evaluation.
@@ 2.1 Instruction Following Tasks
In this section, we specifically focus on multi-turn instruction-following tasks, highlighting their unique characteristics, challenges, and recent benchmark developments. Inspired by task categorization of MT-Bench [21], We group existing multi-turn instruction-following benchmarks into three primary categories: general (mixed) instruction-following, mathematics, and coding. Most benchmarks naturally fall into the general category due to their coverage of diverse task types. Nevertheless, specialized benchmarks targeting math and coding interactions have also emerged prominently, demonstrating unique evaluation dimensions and challenges distinct from general-purpose tasks. Table 1 summarizes these benchmarks, including their evaluation methods, scope, and distinguishing attributes.
@@ 2.1.1 Instruction Following Tasks in General
Recent research has increasingly focused on evaluating the multi-turn interaction abilities of large language models (LLMs), aiming to capture the complexities of real-world dialogue that single-turn benchmarks, such as BIG-Bench [39], CSQA [40], MMLU [41], and GSM8K [42], often fail to address.
@@
MT-Bench [21] emerged as one of the first curated benchmarks designed to evaluate multi-turn instruction following capabilities in LLMs. It consists of 80 two-turn dialogues that span eight categories: writing, roleplay, information extraction, reasoning, math, coding, STEM knowledge, and social science. The benchmark evaluates model responses through pairwise comparisons, assessing correctness and helpfulness. A key contribution of MT-Bench (along with AlpacaEval [43]) is its systematic study of LLM-as-a-judge, where strong LLMs, such as GPT-4, serve as automated evaluators. The study shows that LLM judges achieve over 80% agreement with human evaluators, making them a scalable alternative to human assessments. It also examines potential biases and limitations of LLM judges and suggests mitigation strategies. Since its introduction, LLM-as-a-judge has become a widely adopted evaluation method, shaping benchmarking practices. However, MT-Bench is limited to two-turn dialogues and a relatively small dataset, underscoring the need for more comprehensive benchmarks.
@@
Building on this foundation, MT-Bench++ [24] extends the original MT-Bench dataset by incorporating additional follow-up turns per dialogue, resulting in eight-turn interactions. These extended interactions, enriched with carefully crafted ellipsis and anaphora, further challenge models’ abilities to maintain context and coherence over prolonged exchanges.
@@
Increasing dialogue length alone was insufficient for comprehensive evaluation. MT-Bench-101 [30] introduces a fine-grained taxonomy for multi-turn dialogues. It categorizes interactions into aspects such as perceptivity (context understanding and memory), interactivity (eliciting clarifying questions), and adaptability (reasoning and reflection). With 4,208 turns across 1,388 dialogues and a three-tier ability taxonomy, this benchmark facilitates a detailed assessment of specific interaction skills. Evaluations on 21 prominent LLMs have revealed that even state-of-the-art chat-tuned models exhibit uneven performance across different turns and task types, and that standard alignment techniques (e.g., supervised fine-tuning and RLHF) do not guarantee consistent improvements in multi-turn settings.
@@
MT-Eval [27] takes a different approach by investigating the performance disparities between single-turn and multi-turn interactions. Utilizing 1,170 multi-turn queries derived from human-chat transcripts, MT-Eval categorizes user tasks into four distinct types: follow-up (building upon previous responses), refinement (modifying prior requests), expansion (elaborating on earlier topics), and recollection (retrieving information from earlier turns). The study demonstrates that most LLMs suffer significant performance degradation in multi-turn scenarios, with errors compounding over successive exchanges and the temporal distance from the relevant context further exacerbating the decline This interaction taxonomy has since been adopted by subsequent research, notably M2Lingual [32], which extends multi-turn evaluation into the multilingual domain. M2Lingual applies these interaction categories across 12 diverse languages, revealing concerning cross-lingual brittleness in context retention abilities. The benchmark demonstrates that even advanced models struggle to maintain contextual understanding across language boundaries, with performance deteriorating more severely in non-English interactions, highlighting a critical gap in the multilingual capabilities of current LLMs for sustained dialogues.
@@
M2Lingual is not the only research highlighting such cross-lingual challenges. Multi-IF [36], a benchmark for multi-turn and multilingual instructions, introduces a challenging evaluation set for LLMs involving multi-turn, verifiable writing instructions (e.g., style or format requirements) across 8 languages. The Multi-IF dataset contains 4,501 dialogues created by expanding a single-turn English benchmark (IFEval [12]) into more challenging multi-turn sequences and translating them into 7 additional languages. Concurrently, FairMT-Bench [37] focuses explicitly on fairness in multi-turn dialogues. FairMT-Bench is the first comprehensive benchmark designed to evaluate fairness in open-domain multi-turn dialogues for LLMs, formulating a task taxonomy that targets LLM fairness capabilities across three stages: context understanding, user interaction, and instruction trade-offs—challenges discussed further in Section 4. Based on this taxonomy, the authors constructed fairness datasets, FairMT-1K and FairMT-10K, encompassing two major bias types (stereotype and toxicity) and six bias attributes (age, gender, race, religion, disabled, and appearance), covering nearly all bias categories commonly addressed in fairness evaluation.
@@
While most multi-turn benchmarks focus on context retention and reasoning across turns, FB-Bench [34] introduces a novel dimension by measuring LLMs’ responsiveness to human feedback in multi-turn interaction settings. This Chinese-language benchmark evaluates two crucial aspects of feedback handling: error correction (the ability to fix mistakes when prompted) and response maintenance (the ability to maintain correct responses when challenged). FB-Bench spans diverse task categories, including mathematics, reasoning, coding, text extraction, text error correction, text creation, knowledge Q&A, and text translation. Findings from this benchmark reveal that leading LLMs demonstrate comparable capabilities in error correction across tasks, but their performance varies significantly in response maintenance. Moreover, the research indicates that hinting guidance substantially improves response quality, while exposure to misinformation or fabricated credentials often results in misleading outputs.
@@
Li et al. (2025) [44] recently drew attention to the challenge of maintaining consistent responses in multi-turn LLM interactions, introducing a framework that assumes models should remain “firm” rather than “fickle” when faced with challenging follow-up prompts. Their work introduces the Position-Weighted Consistency (PWC) score, a novel metric that considers the temporal dimension of dialogue by assigning greater penalties to inconsistencies occurring in earlier interaction rounds. This approach reflects the real-world importance of early stability in establishing user trust. Through experiments with leading models across carefully designed challenge scenarios, the authors demonstrated that even high-performing LLMs can be swayed by various follow-up strategies, including emotional appeals and expert authority claims. To enhance response stability, they propose the Confidence-Aware Response Generation (CARG) framework, which integrates model confidence signals into the generation process. Empirical results demonstrate that CARG significantly improves response consistency without compromising accuracy, highlighting the necessity of incorporating turn-based considerations into LLM evaluations.

### v5

@@ Abstract
Recent advances in large language models (LLMs) have substantially improved single-turn task performance, yet real-world applications increasingly demand sophisticated multi-turn interactions. This survey provides a comprehensive review of recent progress in evaluating and enhancing multi-turn LLM interactions. Centered on a task-oriented taxonomy—spanning instruction following in domains such as mathematics and coding, and conversational engagement in role-playing, healthcare, education, and adversarial jailbreak settings—we systematically examine the challenges of maintaining context, coherence, fairness, and responsiveness across prolonged dialogues. We organize existing benchmarks and datasets into coherent categories reflecting the evolving landscape of multi-turn dialogue evaluation, and review a broad spectrum of enhancement methodologies, including model-centric strategies (in-context learning, supervised fine-tuning, reinforcement learning, and architectural innovations), external integration approaches (memory augmentation, retrieval-based methods, and knowledge graphs), and agent-based techniques for collaborative interaction. Finally, we identify open challenges and promising directions for future research to further improve the robustness and effectiveness of multi-turn LLM interactions. Related resources and papers are available at https://github.com/yubol-cmu/Awesome-Multi-Turn-LLMs.
@@ 1 Introduction
The advent of Large Language Models (LLMs), exemplified by influential systems such as GPT series (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023), has significantly reshaped numerous domains, from education and healthcare to customer service and software engineering. These powerful language models demonstrate remarkable proficiency in generating coherent and contextually relevant responses, achieving groundbreaking performances across various language understanding and generation benchmarks.
However, much of the early progress in both the evaluation and improvements of LLMs has been concentrated in single-turn setting, where models are tested on isolated prompts without considering prior conversational context. While this approach has yielded strong benchmark performance, it underexplores LLMs’ broader potential in multi-turn dialogue—the setting that best reflects real-world use. In practice, user needs rarely arise as isolated queries; meaningful and productive interactions typically unfold through sustained, multi-turn exchanges. Effective communication between humans and artificial intelligence inherently demands an understanding of conversational history, nuanced interpretation of previous exchanges, iterative refinement of goals, and adaptive response strategies. Single-turn interaction thus presents a significant limitation, restricting the deployment and utilization of the robust capabilities that LLMs possess.
@@
This emerging field presents rich opportunities and considerable challenges, prompting a rapidly expanding body of research dedicated to optimizing, evaluating, and deploying LLMs within multi-turn settings. Understanding these multi-turn dynamics and systematically addressing their inherent complexities is essential for unlocking the next stage in LLM evolution, thereby significantly expanding their applicability and effectiveness in real-world scenarios.
@@
To provide clarity and facilitate further advancements, this survey specifically addresses multi-turn interaction with LLMs, categorized by tasks along with corresponding evaluation standards, improvement techniques, and current challenges in this domain. We categorize multi-turn interactions according to two main tasks: instruction following tasks and conversational engagement tasks, with the latter encompassing various real-world domain applications. While we refer to several LLM agent-based approaches, we emphasize that LLM-based agents represent a distinct research direction that, though related, falls outside our primary scope. Similarly, despite the growing body of research in Multi-modal LLMs (MLLMs) - including notable benchmarks and frameworks such as ConvBench (Liu et al., 2024b), MMDU (Liu et al., 2024g), MMMT-IF (Epstein et al., 2024), MMDialog Feng et al. (2023), TheaterGen (Cheng et al., 2024), and SVBench (Yang et al., 2025d) - this survey deliberately excludes multimodal capabilities to maintain a focused analysis of text-based multi-turn interactions. Such boundary allows us to provide a comprehensive exploration and in-depth analysis of the landscape of multi-turn interactions specific to LLMs.
@@
Foundational surveys (Zhou et al., 2023; Zeng et al., 2024) provide rigorous analyses of LLMs’ instruction-following ability in single-turn settings. These works assess properties such as faithfulness, robustness, and generalization in static, one-off interactions. While they establish important baselines for instruction alignment, they do not address the dynamic and context-dependent challenges of multi-turn interaction, including cross-turn coherence, state tracking, and adaptation to evolving user intent.
Several surveys on dialogue systems are also closely related. Yi et al. (2025) provides an overview of recent advances in LLM-based multi-turn dialogue systems, while healthcare-oriented dialogue systems are reviewed in (Valizadeh and Parde, 2022; Shi et al., 2024c). In addition, Zhang et al. (2025a) specifically examines LLMs’ multi-turn interaction capabilities, with emphasis on core abilities, evaluation metrics, and algorithmic enhancements for multi-turn contexts.
@@
Instruction-following surveys primarily focus on single-turn settings, and dialogue-system surveys typically emphasize performance within specific domains or tasks. However, the literature lacks a structured, holistic synthesis that compares methods and findings under broad multi-turn interaction settings.
Among existing work, Zhang et al. (2025a) is most closely related to ours. It centers on core multi-turn capabilities, evaluation methodologies, and algorithmic improvements, adopting a largely capability-oriented perspective. In contrast, we take a task-oriented view motivated by real-world deployment, where performance arises from the interplay of multiple capabilities rather than isolated skills. Accordingly, we organize multi-turn interactions by tasks to provide a more application-driven taxonomy. Moreover, Zhang et al. (2025a) offers limited discussion of multi-turn interaction in high-impact settings such as healthcare consultation and educational assistance, where sustained dialogue is essential, and provides comparatively less analysis of improvement strategies and open challenges in multi-turn contexts.
@@
To the best of our knowledge, this survey is the first to provide a comprehensive overview of multi-turn LLM interactions, spanning task formulations, real-world applications, evaluation methodologies, improvement strategies, and key open challenges.
@@ 2 Multi-Turn Interaction Tasks
In this survey, we categorize multi-turn interactions based on tasks rather than capabilities because real-world multi-turn scenarios inherently require integrating multiple capabilities to accomplish a particular objective. Although capabilities such as reasoning, memory, contextual understanding, or adaptability are critical, they rarely function in isolation within conversational contexts. Instead, different capabilities intertwine and complement each other dynamically to achieve specific task goals. By focusing on well defined tasks such as multi turn instruction following and conversational engagement, we better capture the practical complexity of these interactions and provide an intuitive framework for understanding how LLMs satisfy user requirements and perform in real world multi turn conversational settings. Table 1 provides an overview of recent multi-turn instruction-following benchmarks or datasets and their evaluation methods, offering readers a detailed reference for understanding current trends and practices in multi-turn LLM evaluation.
@@ 2.1 Instruction Following Tasks
In this section, we specifically focus on multi-turn instruction-following tasks, highlighting their unique characteristics, challenges, and recent benchmark developments. Inspired by task categorization of MT-Bench (Zheng et al., 2023), We group existing multi-turn instruction-following benchmarks into three primary categories: general (mixed) instruction-following, mathematics, and coding. Most benchmarks naturally fall into the general category due to their coverage of diverse task types. Nevertheless, specialized benchmarks targeting math and coding interactions have also emerged prominently, demonstrating unique evaluation dimensions and challenges distinct from general-purpose tasks.
Dataset Size Data Curation Evaluator Datea Benchmark Dataset Category Total #Dial. Avg. #Turns/Dial. Human-based LLM-based Rule-based Human-as-a-judgeb LLM-as-a-judge Agreement Checkc Evaluation Criteria Mar 2022 MTPB (Nijkamp et al., 2023b) Coding 115 / no no yes no no / Pass rate, Perplexity Dec 2022 ABC-Evald (Finch et al., 2023) Conversation 400 30.3 yes no no yes no / Consistency, Emotion, Understanding, Engagingness, Grammaticality, Informativeness, Quality, Proactivity, Relevance Jun 2023 MathChat-Agent (Wu et al., 2024b) IF-math / 15 no no yes no no / Accuracy Jun 2023 MT-Bench (Zheng et al., 2023) IF-general 80 2 yes no no no yes yes Correctness, Helpfulness Jun 2023 InterCode (Yang et al., 2023) IF-coding / 10 no yes yes no no / Execution Success Sep 2023 MINT (Wang et al., 2024f) IF-reasoning 586 <5<5 no yes yes no no yes Evaluation Quality, Success Rate Oct 2023 MT-Bench++ (Sun et al., 2024) IF-general 80 8 no yes no no yes no Helpfulness, Relevance, Accuracy, Depth, Creativity, Level of Detail Oct 2023 BotChat (Duan et al., 2024) Conversation 547 16 yes yes no no yes / Quality, Similarity to Human Dialogues Jan 2024 EHRAgent (Shi et al., 2024a) IF-coding / / yes no yes no no / Success Rate, Complete Rate Jan 2024 MT-Eval (Kwan et al., 2024) IF-general 168 6.96 no yes yes no yes no Helpfulness, relevance, accuracy, depth, creativity, how well it conforms to constraints; accuracy of classification and recollection Jan 2024 WebLINX ( et al., 2024) IF-general 2337 43 yes no yes no no / Intent Match, Element Similarity Using IoU, Text Similarity Using F1, Turn-Level Score and Overall Score Feb 2024 AQA-Bench (Yang et al., 2025b) IF-reasoning / 15,30e / / yes no no / Goal Metric, Policy Metric (Both for Each Interactive Environment) Feb 2024 MT-Bench-101 (Bai et al., 2024) IF-general 1388 3.03 no yes no no yes yes Perceptivity (Context Memory, Context Understanding, Context Interference), Adaptability (Rephrasing, Reflection, Reasoning), Interactivity (Questioning) May 2024 MathChat-Bench (Liang et al., 2024) IF-math 5276 4 no yes no no yes yes Follow-Up QA, Error Correction, Error Analysis, Problem Generation Jun 2024 M2Lingualf (Maheshwary et al., 2025) IF-general 1140 2.48 no yes MT-Bench same as MT-Bench’s Aug 2024 SysBench (Qin et al., 2025) IF-general 500 5 yes no yes no no / Constraint Satisfaction Rate, Instruction Satisfaction Rate, Session Stability Rate Oct 2024 FB-Bench (Li et al., 2025b) IF-general 591 2 yes yes no no yes yes Error Correction, Response Maintenance Oct 2024 WILT (Banatt et al., 2024) IF-reasoning 50 <30<30 / / yes no no / Accuracy in Deducing the Hidden Functions Oct 2024 Multi-IF (He et al., 2024) IF-general 909 3 no yes yes no no / Instruction- and Conversation-Level Accuracy of Instruction Following Oct 2024 FairMT-Bench (Fan et al., 2025) IF-fairness 10k 4 no yes no no yes yes Direct and Implicit Bias Dec 2024 MMSQL (Guo et al., 2025) IF-coding 6493 6 no yes yes no yes yes Exact Matching, Execution Accuracy, Dual Assessment of Question Type Detection, Execution Accuracy, Response Quality Score
  • a
    We use the date of their initial version on arXiv to rank them, allowing the development trajectory
    and patterns to be revealed.
  • b
    Do not include the human evaluation done for agreement check between LLM and human.
  • c
    If LLM-as-a-judge is used, whether there is an agreement check between LLM and human
    annotators on a subset of the dataset.
  • d
    We intentionally include ABC-Eval and BotChat in this table as they represent early explorations
    in multi-turn interactions. Their presence illustrates the developmental trajectory of the field,
    although a more detailed discussion can be found in §2.2.
  • e
    15 for easy mode, 30 for hard mode.
  • f
    M2Lingual includes dialogues for 70+ languages. The statistics shown here are restricted to only
    English multi-turn dialogues.
Table 1: An Overview of Recent Multi-Turn Instruction Following Benchmark or Datasets and Evaluation Methods.
@@ 2.1.1 Instruction Following Tasks in General
Recent research has increasingly focused on evaluating the multi-turn interaction abilities of large language models (LLMs), aiming to capture the complexities of real-world dialogue that single-turn benchmarks, such as BIG-Bench (Ghazal et al., 2013), CSQA (Talmor et al., 2019), MMLU (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021), often fail to address.
@@
MT-Bench (Zheng et al., 2023) emerged as one of the first curated benchmarks designed to evaluate multi-turn instruction following capabilities in LLMs. It consists of 80 two-turn dialogues that span eight categories: writing, roleplay, information extraction, reasoning, math, coding, STEM knowledge, and social science. The benchmark evaluates model responses through pairwise comparisons, assessing correctness and helpfulness. A key contribution of MT-Bench (along with AlpacaEval (Li et al., 2023e)) is its systematic study of LLM-as-a-judge, where strong LLMs, such as GPT-4, serve as automated evaluators. The study shows that LLM judges achieve over 80% agreement with human evaluators, making them a scalable alternative to human assessments. It also examines potential biases and limitations of LLM judges and suggests mitigation strategies. Since its introduction, LLM-as-a-judge has become a widely adopted evaluation method, shaping benchmarking practices. However, MT-Bench is limited to two-turn dialogues and a relatively small dataset, underscoring the need for more comprehensive benchmarks.
@@
Building on this foundation, MT-Bench++ (Sun et al., 2024) extends the original MT-Bench dataset by incorporating additional follow-up turns per dialogue, resulting in eight-turn interactions. These extended interactions, enriched with carefully crafted ellipsis and anaphora, further challenge models’ abilities to maintain context and coherence over prolonged exchanges.
@@
Increasing dialogue length alone was insufficient for comprehensive evaluation. MT-Bench-101 (Bai et al., 2024) introduces a fine-grained taxonomy for multi-turn dialogues. It categorizes interactions into aspects such as perceptivity (context understanding and memory), interactivity (eliciting clarifying questions), and adaptability (reasoning and reflection). With 4,208 turns across 1,388 dialogues and a three-tier ability taxonomy, this benchmark facilitates a detailed assessment of specific interaction skills. Evaluations on 21 prominent LLMs have revealed that even state-of-the-art chat-tuned models exhibit uneven performance across different turns and task types, and that standard alignment techniques (e.g., supervised fine-tuning and RLHF) do not guarantee consistent improvements in multi-turn settings.
@@
MT-Eval (Kwan et al., 2024) takes a different approach by investigating the performance disparities between single-turn and multi-turn interactions. Utilizing 1,170 multi-turn queries derived from human-chat transcripts, MT-Eval categorizes user tasks into four distinct types: follow-up (building upon previous responses), refinement (modifying prior requests), expansion (elaborating on earlier topics), and recollection (retrieving information from earlier turns). The study demonstrates that most LLMs suffer significant performance degradation in multi-turn scenarios, with errors compounding over successive exchanges and the temporal distance from the relevant context further exacerbating the decline This interaction taxonomy has since been adopted by subsequent research, notably M2Lingual (Maheshwary et al., 2025), which extends multi-turn evaluation into the multilingual domain. M2Lingual applies these interaction categories across 12 diverse languages, revealing concerning cross-lingual brittleness in context retention abilities. The benchmark demonstrates that even advanced models struggle to maintain contextual understanding across language boundaries, with performance deteriorating more severely in non-English interactions, highlighting a critical gap in the multilingual capabilities of current LLMs for sustained dialogues.
@@
M2Lingual is not the only research highlighting such cross-lingual challenges. Multi-IF (He et al., 2024), a benchmark for multi-turn and multilingual instructions, introduces a challenging evaluation set for LLMs involving multi-turn, verifiable writing instructions (e.g., style or format requirements) across 8 languages. The Multi-IF dataset contains 4,501 dialogues created by expanding a single-turn English benchmark (IFEval (Zhou et al., 2023)) into more challenging multi-turn sequences and translating them into 7 additional languages. Concurrently, FairMT-Bench (Fan et al., 2025) focuses explicitly on fairness in multi-turn dialogues. FairMT-Bench is the first comprehensive benchmark designed to evaluate fairness in open-domain multi-turn dialogues for LLMs, formulating a task taxonomy that targets LLM fairness capabilities across three stages: context understanding, user interaction, and instruction trade-offs—challenges discussed further in Section 4. Based on this taxonomy, the authors constructed fairness datasets, FairMT-1K and FairMT-10K, encompassing two major bias types (stereotype and toxicity) and six bias attributes (age, gender, race, religion, disabled, and appearance), covering nearly all bias categories commonly addressed in fairness evaluation.
@@
While most multi-turn benchmarks focus on context retention and reasoning across turns, FB-Bench (Li et al., 2025b) introduces a novel dimension by measuring LLMs’ responsiveness to human feedback in multi-turn interaction settings. This Chinese-language benchmark evaluates two crucial aspects of feedback handling: error correction (the ability to fix mistakes when prompted) and response maintenance (the ability to maintain correct responses when challenged). FB-Bench spans diverse task categories, including mathematics, reasoning, coding, text extraction, text error correction, text creation, knowledge Q&A, and text translation. Findings from this benchmark reveal that leading LLMs demonstrate comparable capabilities in error correction across tasks, but their performance varies significantly in response maintenance. Moreover, the research indicates that hinting guidance substantially improves response quality, while exposure to misinformation or fabricated credentials often results in misleading outputs.
@@
Li et al. (2025c) recently draw attention to the challenge of maintaining consistent responses in multi-turn LLM interactions, introducing a framework that assumes models should remain “firm” rather than “fickle” when faced with challenging follow-up prompts. Their work introduces the Position-Weighted Consistency (PWC) score, a novel metric that considers the temporal dimension of dialogue by assigning greater penalties to inconsistencies occurring in earlier interaction rounds. This approach reflects the real-world importance of early stability in establishing user trust. Through experiments with leading models across carefully designed challenge scenarios, the authors demonstrated that even high-performing LLMs can be swayed by various follow-up strategies, including emotional appeals and expert authority claims. To enhance response stability, they propose the Confidence-Aware Response Generation (CARG) framework, which integrates model confidence signals into the generation process. Empirical results demonstrate that CARG significantly improves response consistency without compromising accuracy, highlighting the necessity of incorporating turn-based considerations into LLM evaluations.